# Impact of Systematic Redshift Errors on the Cross-correlation of the Lyman-$\alpha$ Forest with Quasars at Small Scales Using DESI Early Data

### v1

@@ 1 Introduction
[…]
@@ 2.1 Quasar Sample
(a) Redshift distribution for the tracer quasar sample (without BALs) used in this paper. The red hatched histogram represents the Lyman-$\alpha$ forest sample. The dashed grey line represents the lower bound on the tracer quasar sample, $z=1.88$.
(b) Redshift distribution of Lyman-$\alpha$ forest pixels output during the continuum fitting process for the baseline analysis.
Figure 1: Redshift and Lyman-$\alpha$ forest pixel distributions for the tracer quasar and Lyman-$\alpha$ forest samples used in this work.
@@
After these cuts, a total of 290,506 quasars are present in the quasar catalog. However, only the 106,861 quasars with redshift z > 1.88 will be included in the cross-correlation discussed in section 3.2. We refer to these quasars as tracer quasars or the tracer quasar catalog. Figure 0(a) shows the redshift distribution of the entire quasar sample with the dashed grey line representing this cutoff for the tracer quasar sample. The red hatched histogram shows the redshift distribution for the Lyman-$\alpha$ sample discussed in the next section.
@@ 2.2 Lyman-$\alpha$ Forest Sample
With these redshift cuts, there are 109,900 valid Lyman-$\alpha$ forests from the quasar catalog. Prior to fitting the continuum (described in section 3.1), we first mask regions of the spectra that are contaminated. This contamination can come from sky lines, galactic absorption, BALs, and Damped Lyman Alpha (DLA) regions. We mask any DLA regions identified with the DLAFinder [36] and BAL regions identified with the BALFinder [34] of the quasar spectra. For the sky lines and galactic absorption, we mask four areas of the spectra: sky lines from [5570.5, 5586.7] Å; and the Ca H, and K lines from the Milky Way from [3967.3, 3971.0] Å and [3933.0, 3935.8] Å, respectively. We also perform a calibration using the CIII between 1600 < $\lambda_{\text{RF}}$ [Å] < 1850, which is shaded in orange in figure 2. This calibration step follows the same process described in section 3.1 and in [35]. We require each forest to have a minimum length of 40 Å. Forests that are too short (don’t have enough pixels), have low signal (< 0), or have problems fitting the continuum (negative continuum) will be rejected during the calibration and continuum fitting process (section 3.1). After these rejections, the final Lyman-$\alpha$ sample has 88,511 forests. Figure 0(b) shows the redshift distribution of Lyman-$\alpha$ pixels in the sample.
@@ 3.3 The Distortion and Covariance Matrices
We follow the same method as in [14] to estimate the covariance matrix for the cross-correlation. The covariance matrix is estimated by splitting the sky into $s$ regions (according to HEALPix pixels) and then computing their weighted covariance. Assuming that correlations between regions are negligible, the covariance is written as
@@
The covariance matrix is dominated by the diagonal elements (the variance) of the matrix. The estimates of off-diagonal elements are noisy, so they are modeled as a function of the difference of separation along the line of sight ($\Delta r_{\parallel}=|r^{\text{A}}_{\parallel}-r^{\text{B}}_{\parallel}|$) and across the line of sight ($\Delta r_{\perp}=|r^{\text{A}}_{\perp}-r^{\text{B}}_{\perp}|$) and then smoothed [8]. Correlation coefficients that have the same $\Delta r_{\parallel},\Delta r_{\perp}$ are averaged.
@@ 3.4 The Model of the Lyman-$\alpha$ Forest-Quasar Cross-correlation
This section presents the theoretical model for the auto- and cross-correlations. Though this work focuses on the cross-correlation, the auto-correlation is necessary to include during the fitting process as it helps break degeneracies between parameters.
@@ 4 Validation with Mock Data
To ensure that the choice of bin width does not introduce any biases to our analysis, and especially to the two redshift error parameters $\Delta r_{\parallel}$ and $\sigma_{\text{v}}$, we run our tests using mock data for three different cross-correlation bin widths: $1\,h^{-1}\,\mathrm{Mpc}$, $2\,h^{-1}\,\mathrm{Mpc}$, and $4\,h^{-1}\,\mathrm{Mpc}$. We follow the same method described in section 3, however we do not need to perform the calibration step on mock data (section 3.1). For the cross-correlation, we still restrict $r_{\parallel}$, $r_{\perp}$in $[-80,80]\,h^{-1}\,\mathrm{Mpc}$. For the $2\,h^{-1}\,\mathrm{Mpc}$ bin width, this gives 80 bins along the line of sight, 40 bins across the line of sight and N = 3,200 bins in total. For the $4\,h^{-1}\,\mathrm{Mpc}$ bin width, this gives 40 bins along the line of sight, 20 bins across the line of sight, and N = 800 bins in total. Since the auto-correlation doesn’t help to constrain redshift errors we choose to follow the setup from [14] as we use the auto-correlations from [14] in section 5. The free and fixed parameters are described in table 5 in appendix A.
@@
The mocks studied here do not have any systematic errors added to the quasar redshift so we expect $\Delta r_{\parallel}$ $\sim 0$. The true value of $\sigma_{\text{v}}$ is known and we do not expect it to be 0, as the mocks include non-linear effects. The results for these parameters for each mock and each bin width tested are shown in figure 4. The weighted average across all mocks is shown by the hatched shaded regions. For the $1\,h^{-1}\,\mathrm{Mpc}$ bin width: $\Delta r_{\parallel}$ = $0.039\pm 0.021$ $h^{-1}$ Mpc and $\sigma_{\text{v}}=4.64\pm 0.11$ $h^{-1}$ Mpc. For the $2\,h^{-1}\,\mathrm{Mpc}$ bin width: $\Delta r_{\parallel}$ = $0.079\pm 0.030$ $h^{-1}$ Mpc and $\sigma_{\text{v}}=4.05\pm 0.11$ $h^{-1}$ Mpc. For the $4\,h^{-1}\,\mathrm{Mpc}$ bin width: $\Delta r_{\parallel}$ = $0.10\pm 0.04$ $h^{-1}$ Mpc and $\sigma_{\text{v}}=3.88\pm 0.11$ $h^{-1}$ Mpc. The $\chi^{2}$ values averaged over all mocks are 1.023 for the $1\,h^{-1}\,\mathrm{Mpc}$ bin width, 1.022 for the $2\,h^{-1}\,\mathrm{Mpc}$ bin width, and 1.009 for the $4\,h^{-1}\,\mathrm{Mpc}$ bin width, which all indicate a good fit.
@@ 5.1 Cross-correlation and Baseline Fit Results
Table 2: Results of the fits for the baseline, BAL, and the ZBAL configurations. The number of quasars above $z=1.88$ in each catalog is given. In all fits the BAO parameters are fixed at 1.0. The rest of the parameters are allowed to float in each fit and are described in 5 in appendix A. Each fit has 11,644 data bins, 13 free parameters and we also list the effective redshift, $z_{\text{eff}}$. Since each fit is a joint fit of the auto- and cross-correlations, we list the $\chi^{2}$ value given for the joint fit.
@@
The apparent underestimation of DESI redshifts persisted when updating the quasar templates in Redrock for processing the DESI year one quasar sample [27]. In section 6, we discuss the improvements in redshift performance for DESI year 1 data at $z>2$ when using an updated version of Redrock and quasar templates that properly model Lyman-$\alpha$ optical depth.
@@ 5.2 Dependence on Quasar Redshift
$A_{\text{inst}}$ $2.29\pm 0.29$ $2.38\pm 0.27$ $2.42\pm 0.27$ $2.42\pm 0.27$
$\chi^{2}_{\text{ JOINT}}$ 1.033 1.030 1.030 1.053
@@
Table 3: Results of the fits for the four redshift evolution bins. The number of quasars in each bin is given, along with the redshift range of each bin. In all fits the BAO parameters are fixed at 1.0. The rest of the parameters are allowed to float in each fit and are described in table 5 in appendix A. Each fit has 11,644 data bins, 13 free parameters and we also list the effective redshift, $z_{\text{eff}}$. Since each fit is a joint fit of the auto- and cross-correlations, we list the $\chi^{2}$ value given for the joint fit.
@@ 5.3 Impact of BALs and Updating BAL Quasar Redshifts
Figure 7: Results on $\Delta r_{\parallel}$ (left, blue points) and $\sigma_{\text{v}}$ (right, blue triangles) for the four different configurations presented in this work: baseline, BAL, and ZBAL. The shaded area represents the value of the parameters from the baseline fit.
[…]
@@
To test if updating the BAL quasar redshifts after the BAL regions are masked improves the $\Delta r_{\parallel}$ parameter, we use the quasar catalog from [19]. This is the same catalog as described above, but with updated redshifts for the BAL quasars, of which 28,185 BAL quasars have an updated redshift. Because the redshifts are updated, we can no longer use the Lyman-$\alpha$ forest sample described in [35]. We follow the method in [35] and in section 3 for the continuum fitting process and our final Lyman-$\alpha$ sample consists of 88,432 forests. We use the tracer catalog with updated redshifts and follow the method described in section 3 to measure and fit the correlations. We can no longer use the auto-correlation from [14], so we measure those following the method in [14]. The results are shown in figure 7 and in table 2 in the ZBAL column. We find $\Delta r_{\parallel}$ = $-1.90\pm 0.11$  $h^{-1}$ Mpc and $\sigma_{\text{v}}=7.0\pm 0.6$  $h^{-1}$ Mpc. The value of $\Delta r_{\parallel}$ has shifted in the negative direction and is now equal ($0.35\sigma$) to the baseline value. This negative shift is consistent with what is found in [19], which found that the redshifts of BAL quasars after masking were shifted to the blue.
This shows that BAL quasars can affect the correlations when they are included in the tracer catalog and their redshifts are not updated after the BAL regions are masked. When the redshifts are updated the inclusion of BAL quasars in the tracer catalog has no effect.
@@ 6 Addressing the Redshift Dependency
is the effective optical depth of the Lyman-$\alpha$ transition, $\tau_{0}=(5.54\pm 0.64)\times 10^{-3}$ and $\gamma=3.182\pm 0.074$ [61]. $\overline{F}$ depends on $\tau_{\text{eff},\alpha}$ at $z=\frac{\lambda_{\text{obs}}-\lambda_{\text{Ly}\alpha}}{\lambda_{\text{Ly}\alpha}}$ according to eq. 6.1. $\tau_{\text{eff},\alpha}$ evolves with redshift, as parameterized by eq. 6.2 [e.g., 62, 63, 61], implying greater overall flux suppression in the Lyman-$\alpha$ forest region of quasar spectra at higher redshifts. The BOSS quasar templates used for EDR and the DESI year 1 quasar templates were derived from samples of quasar spectra that were not corrected for the optical depth of Lyman-$\alpha$ photons. On the other hand, a real-time correction for this flux suppression was being applied in Redrock during redshift fitting, leading to an over-suppression of template flux relative to what was expected at $\lambda_{\text{RF}}<1215.67$Å. This resulted in a bias that increases with the amount of suppression (i.e. redshift) at $z>2$.
@@
[…]
@@
Parameter baseline z1 z2 z3 z4 No. quasars 510,501 201,290 138,364 82,710 88,137 Redshift range - $z\leq 2.2$ $2.2<z\leq 2.5$ $2.5<z\leq 2.8$ $z>2.8$ $b_{\text{Ly}\alpha}$ $-0.0543\pm 0.0035$ $-0.049\pm 0.005$ $-0.055\pm 0.005$ $-0.060\pm 0.006$ $-0.162\pm 0.020$ $\beta_{\text{Ly}\alpha}$ $3.78\pm 0.26$ $3.3\pm 0.4$ $3.8\pm 0.4$ $4.5\pm 0.5$ $2.01\pm 0.28$ $10^{3}b_{\eta,\text{SiII(1190)}}$ $-2.5\pm 0.4$ $-2.6\pm 0.4$ $-2.8\pm 0.4$ $-3.9\pm 0.5$ $-3.4\pm 0.5$ $10^{3}b_{\eta,\text{SiII(1193)}}$ $-1.34\pm 0.35$ $-1.3\pm 0.4$ $-2.2\pm 0.4$ $-2.0\pm 0.5$ $-2.2\pm 0.5$ $10^{3}b_{\eta,\text{SiIII(1207)}}$ $-4.50\pm 0.31$ $-5.0\pm 0.4$ $-5.0\pm 0.4$ $-6.0\pm 0.5$ $-5.4\pm 0.5$ $10^{3}b_{\eta,\text{SiII(1260)}}$ $-3.1\pm 0.6$ $-2.9\pm 0.5$ $-3.2\pm 0.6$ $-3.8\pm 0.7$ $-3.2\pm 0.6$ $b_{\text{HCD}}$ $-0.0574\pm 0.0034$ $-0.0491\pm 0.0029$ $-0.061\pm 0.004$ $-0.082\pm 0.005$ $-0.089\pm 0.006$ $b_{\text{q}}$ $3.97\pm 0.10$ $3.35\pm 0.13$ $3.99\pm 0.13$ $4.73\pm 0.17$ $4.35\pm 0.23$ $\Delta r_{\parallel}$ ($h^{-1}$pc) $-0.08\pm 0.04$ $-0.01\pm 0.06$ $-0.16\pm 0.07$ $-0.09\pm 0.08$ $-0.11\pm 0.13$ $\sigma_{\text{v}}$ ($h^{-1}$pc) $5.21\pm 0.17$ $4.79\pm 0.25$ $5.35\pm 0.25$ $5.45\pm 0.35$ $3.46\pm 0.35$ $\xi_{0}^{\text{TP}}$ $0.074\pm 0.008$ $0.066\pm 0.013$ $0.061\pm 0.014$ $0.114\pm 0.023$ $0.52\pm 0.09$ $10^{4}A_{\text{inst}}$ $1.33\pm 0.09$ $1.4\pm 0.1$ $1.36\pm 0.10$ $1.35\pm 0.10$ $1.51\pm 0.10$ $\chi^{2}_{\text{ JOINT}}$ 1.027 1.025 1.022 0.999 0.995 $z_{\text{eff}}$ 2.32 2.09 2.33 2.61 2.98
Table 4: Results of the fits using the updated quasar templates for the new baseline and four redshift bins. The number of quasars in each bin is given, along with the redshift range of each bin where applicable. The free parameters are the same as described in section 5. In all fits the BAO parameters are fixed at 1.0. Each fit has 11,644 data bins, 13 free parameters and we also list the effective redshift, $z_{\text{eff}}$. Since each fit is a joint fit of the auto- and cross-correlations, we list the $\chi^{2}$ value given for the joint fit.

### v2

@@ 1 Introduction
In order to get the high precision needed for the cosmological measurements from BAO with the Lyman-$\alpha$ forest-quasar cross-correlation, the redshifts that DESI measures for each quasar spectrum need to be extremely precise. While the redshift measurement of the majority of DESI galaxies, which rely on the presence of narrow emission lines and spectral features, is indeed very precise, this is not the case for quasars whose emission lines are essentially broad, and complex. There are therefore many factors that can cause errors on redshift measurements. Quasars that have Broad Absorption Line (BAL) features tend to have larger errors on redshift since the BAL features tend to affect the absorption blueward of the emission line centers. The resulting emission line asymmetry can shift the redshift estimate [18, 19]. Additionally, quasar emission line profiles can vary from quasar to quasar, and can be shifted from the systemic redshift. This is especially true for high-ionization, broad emission lines like CIV, which can be blueshifted a few hundred km s-1 from low-ionization, broad emission lines like MgII which are often closer to the systemic redshift [20, 21]. Emission lines can also be broadened, and even asymmetric, for example Lyman-$\alpha$ absorption in the IGM can cause an asymmetry in the line profile [22, 20].
@@ 2.1 Quasar Sample
Figure 1: The left plot shows the redshift distribution for the tracer quasar sample (without BALs) used in this paper. The red hatched histogram represents the Lyman-$\alpha$ forest sample. The dashed grey line represents the lower bound on the tracer quasar sample, $z=1.88$. The right plot shows the redshift distribution of the Lyman-$\alpha$ forest pixels output during the continuum fitting process for the baseline analysis.
@@
After these cuts, a total of 290,506 quasars are present in the quasar catalog. However, only the 106,861 quasars with redshift z > 1.88 will be included in the cross-correlation discussed in section 3.2. We refer to these quasars as tracer quasars or the tracer quasar catalog. Figure 1 shows the redshift distribution of the entire quasar sample with the dashed grey line representing this cutoff for the tracer quasar sample. The red hatched histogram shows the redshift distribution for the Lyman-$\alpha$ sample discussed in the next section.
@@ 2.2 Lyman-$\alpha$ Forest Sample
With these redshift cuts, there are 109,900 valid Lyman-$\alpha$ forests from the quasar catalog. Prior to fitting the continuum (described in section 3.1), we first mask regions of the spectra that are contaminated. This contamination can come from sky lines, galactic absorption, BALs, and Damped Lyman Alpha (DLA) regions. We mask any DLA regions identified with the DLAFinder [36] and BAL regions identified with the BALFinder [34] of the quasar spectra. For the sky lines and galactic absorption, we mask four areas of the spectra: sky lines from [5570.5, 5586.7] Å; and the Ca H, and K lines from the Milky Way from [3967.3, 3971.0] Å and [3933.0, 3935.8] Å, respectively. We also perform a calibration using the CIII between 1600 < $\lambda_{\text{RF}}$ [Å] < 1850, which is shaded in orange in figure 2. This calibration step follows the same process described in section 3.1 and in [35]. We require each forest to have a minimum length of 40 Å. Forests that are too short (don’t have enough pixels), have low signal (< 0), or have problems fitting the continuum (negative continuum) will be rejected during the calibration and continuum fitting process (section 3.1). After these rejections, the final Lyman-$\alpha$ sample has 88,511 forests. Figure 1 shows the redshift distribution of Lyman-$\alpha$ pixels in the sample.
@@ 3.3 The Distortion and Covariance Matrices
We follow the same method as in [14] to estimate the covariance matrix for the cross-correlation. The covariance matrix is estimated by splitting the sky into $s$ regions (according to HEALPix [40] pixels) and then computing their weighted covariance. For our analysis we use 440 HEALPix pixels with nside = 16. Assuming that correlations between regions are negligible, the covariance is written as
@@
The cross-covariance matrix has N = $12800\times 12800=163$ million entries and the auto-covariance (from [14]) has N = $2500\times 2500=6.25$ million entries. The covariance matrices are dominated by the diagonal elements of the matrix. The estimates of off-diagonal elements are noisy, so they are modeled as a function of the difference of separation along the line of sight ($\Delta r_{\parallel}=|r^{\text{A}}_{\parallel}-r^{\text{B}}_{\parallel}|$) and across the line of sight ($\Delta r_{\perp}=|r^{\text{A}}_{\perp}-r^{\text{B}}_{\perp}|$) and then smoothed [8]. This smoothing is necessary to obtain a positive-definite, or invertible, covariance matrix. Correlation coefficients that have the same $\Delta r_{\parallel},\Delta r_{\perp}$ are averaged.
@@ 3.4 The Model of the Lyman-$\alpha$ Forest-Quasar Cross-correlation
This section presents the theoretical model for the cross-correlation. Though this work focuses on the cross-correlation, the auto-correlation is necessary to include during the fitting process as it helps break degeneracies between parameters. We follow the method from [14] for fitting the auto-correlation, and use the auto-correlations from [14] in the baseline analysis.
@@ 4 Validation with Mock Data
To ensure that the choice of bin width does not introduce any biases to our analysis, and especially to the two redshift error parameters $\Delta r_{\parallel}$ and $\sigma_{\text{v}}$, we run our tests using mock data for three different cross-correlation bin widths: $1\,h^{-1}\,\mathrm{Mpc}$, $2\,h^{-1}\,\mathrm{Mpc}$, and $4\,h^{-1}\,\mathrm{Mpc}$. We follow the same method described in section 3, however we do not need to perform the calibration step on mock data (section 3.1) during the continuum fitting process. For the cross-correlation, we still restrict $r_{\parallel}$, $r_{\perp}$in $[-80,80]\,h^{-1}\,\mathrm{Mpc}$. For the $2\,h^{-1}\,\mathrm{Mpc}$ bin width, this gives 80 bins along the line of sight, 40 bins across the line of sight and N = 3,200 bins in total. For the $4\,h^{-1}\,\mathrm{Mpc}$ bin width, this gives 40 bins along the line of sight, 20 bins across the line of sight, and N = 800 bins in total. Though the auto-correlation does not help to constrain redshift errors we include it to help break degeneracies with other parameters. We follow the setup and method from [14] for the auto-correlation. We restrict $r_{\parallel}$, $r_{\perp}$ $\in[0,200]$ $h^{-1}$ Mpc using 50 bins with a 4  $h^{-1}$ Mpc bin width, which gives 2500 total bins. The free and fixed parameters are described in table 5 in appendix A.
@@
The mocks studied here do not have any systematic errors added to the quasar redshift so we expect $\Delta r_{\parallel}$ $\sim 0$. The true value of $\sigma_{\text{v}}$ is known and we do not expect it to be 0, as the mocks include non-linear effects. The results for these parameters for each mock and each bin width tested are shown in figure 4. The weighted average across all mocks is shown by the hatched shaded regions. For the $1\,h^{-1}\,\mathrm{Mpc}$ bin width: $\Delta r_{\parallel}$ = $0.039\pm 0.021$ $h^{-1}$ Mpc and $\sigma_{\text{v}}=4.64\pm 0.11$ $h^{-1}$ Mpc. For the $2\,h^{-1}\,\mathrm{Mpc}$ bin width: $\Delta r_{\parallel}$ = $0.079\pm 0.030$ $h^{-1}$ Mpc and $\sigma_{\text{v}}=4.05\pm 0.11$ $h^{-1}$ Mpc. For the $4\,h^{-1}\,\mathrm{Mpc}$ bin width: $\Delta r_{\parallel}$ = $0.10\pm 0.04$ $h^{-1}$ Mpc and $\sigma_{\text{v}}=3.88\pm 0.11$ $h^{-1}$ Mpc. The $\chi^{2}$ values averaged over all mocks are $11908/(11644-12)=1.023$ for the $1\,h^{-1}\,\mathrm{Mpc}$ bin width, $4182/(4102-12)=1.022$ for the $2\,h^{-1}\,\mathrm{Mpc}$ bin width, and $2230/(2222-12)=1.009$ for the $4\,h^{-1}\,\mathrm{Mpc}$ bin width, which all indicate a good fit.
@@ 5.1 Cross-correlation and Baseline Fit Results
Table 2: Results of the fits for the baseline, BAL, and the ZBAL configurations. The number of quasars above $z=1.88$ in each catalog is given. In all fits the BAO parameters are fixed at 1.0. The rest of the parameters are allowed to float in each fit and are described in Table 5 in appendix A. Each fit has 11,644 data bins, 13 free parameters and we also list the effective redshift, $z_{\text{eff}}$. Since each fit is a joint fit of the auto- and cross-correlations, we list the $\chi^{2}$ and reduced $\chi^{2}$ ($\chi^{2}$/dof) values given for the joint fit.
@@
The apparent underestimation of DESI redshifts persisted when updating the quasar templates in Redrock for processing the DESI year one quasar sample, as they did not account for the Lyman-$\alpha$ optical depth [27]. In section 6, we discuss the improvements in redshift performance for DESI year 1 data at $z>2$ when using an updated version of Redrock and quasar templates that properly model Lyman-$\alpha$ optical depth.
@@ 5.2 Dependence on Quasar Redshift
$10^{4}A_{\text{inst}}$ $2.29\pm 0.29$ $2.38\pm 0.27$ $2.42\pm 0.27$ $2.42\pm 0.27$
$\chi^{2}$ 12010.9 11978.3 11980.5 12248.9
$\chi^{2}$/dof 1.033 1.030 1.030 1.053
@@
Table 3: Results of the fits for the four redshift evolution bins. The number of quasars in each bin is given, along with the redshift range of each bin. In all fits the BAO parameters are fixed at 1.0. The rest of the parameters are allowed to float in each fit and are described in table 5 in appendix A. Each fit has 11,644 data bins, 13 free parameters and we also list the effective redshift, $z_{\text{eff}}$. Since each fit is a joint fit of the auto- and cross-correlations, we list the $\chi^{2}$ and reduced $\chi^{2}$ ($\chi^{2}$/dof) values given for the joint fit.
@@ 5.3 Impact of BALs and Updating BAL Quasar Redshifts
Figure 7: Results on $\Delta r_{\parallel}$ (left, blue points) and $\sigma_{\text{v}}$ (right, blue triangles) for the three different configurations presented in this work: baseline, BAL, and ZBAL. The shaded area represents the value of the parameters from the baseline fit.
BAL systems can impact the shape of quasar emission lines and introduce shifts in the estimated quasar redshift, up to several hundred km s-1 [19]. BAL quasars often have larger redshift errors compared to quasars with no BAL features because the BAL features impact the blue side of emission lines [18]. This is true in DESI, where Redrock tends to overestimate redshifts for BAL quasars [18]. The current pipeline used for most Lyman-$\alpha$ studies, Picca, has the option to mask out BAL regions during the continuum fitting process, but does not update the quasar redshift when these masks are applied so any systematic errors on the quasar redshift will still be present. [19] shows that when BAL regions of a quasar spectrum are masked and new redshifts are obtained, those new redshifts shift by 240 km s-1 on average. We explore in this section the effect of including the BAL quasars in the tracer catalog as well as using the updated BAL redshifts from [19] once the BAL regions of the quasar spectra are masked.
@@
To test if updating the BAL quasar redshifts after the BAL regions are masked improves the $\Delta r_{\parallel}$ parameter, we use the quasar catalog from [19]. This is the same catalog as described above, but with updated redshifts for the BAL quasars, of which 28,185 BAL quasars have an updated redshift. Because the redshifts are updated, the Lyman-$\alpha$ forest sample changes slightly, both in the number of forests and the redshift distribution of pixels. While the effect is small, we make the choice to not use the Lyman-$\alpha$ forest sample described in [35]. We follow the method in [35] and in section 3 for the continuum fitting process and our final Lyman-$\alpha$ sample consists of 88,432 forests. We use the tracer catalog with updated redshifts and follow the method described in section 3 to measure and fit the correlations. We can no longer use the auto-correlation from [14], so we measure those following the method in [14]. The results are shown in figure 7 and in table 2 in the ZBAL column. We find $\Delta r_{\parallel}$ = $-1.90\pm 0.11$  $h^{-1}$ Mpc and $\sigma_{\text{v}}=7.0\pm 0.6$  $h^{-1}$ Mpc. The value of $\Delta r_{\parallel}$ has shifted in the negative direction and is now equal ($0.35\sigma$) to the baseline value. This negative shift is consistent with what is found in [19], which found that the redshifts of BAL quasars after masking were shifted to the blue.
In both configurations when BAL quasars are added in at the catalog level the values of $b_{\text{Ly}\alpha}$ and $\beta_{\text{Ly}\alpha}$ change by a factor of two compared to the baseline. These parameters are highly correlated with each other and with $b_{\text{q}}$ ($\rho\sim 0.9$). By adding the BAL quasars at the catalog level we are changing the value of $b_{\text{q}}$, so it is not unexpected that $b_{\text{Ly}\alpha}$ and $\beta_{\text{Ly}\alpha}$ also change.
This shows that BAL quasars can affect the correlations when they are included in the tracer catalog and their redshifts are not updated after the BAL regions are masked. When the redshifts are updated the inclusion of BAL quasars in the tracer catalog has no effect on $\Delta r_{\parallel}$.
@@ 6 Addressing the Redshift Dependency
is the effective optical depth of the Lyman-$\alpha$ transition, $\tau_{0}=(5.54\pm 0.64)\times 10^{-3}$ and $\gamma=3.182\pm 0.074$ [62]. $\overline{F}$ depends on $\tau_{\text{eff},\alpha}$ at $z=\frac{\lambda_{\text{obs}}-\lambda_{\text{Ly}\alpha}}{\lambda_{\text{Ly}\alpha}}$ according to eq. 6.1. $\tau_{\text{eff},\alpha}$ evolves with redshift, as parameterized by eq. 6.2 [e.g., 63, 64, 62], implying greater overall flux suppression in the Lyman-$\alpha$ forest region of quasar spectra at higher redshifts. The BOSS quasar templates used for EDR and the DESI year 1 quasar templates were derived from samples of quasar spectra that were not corrected for the optical depth of Lyman-$\alpha$ photons. On the other hand, a real-time correction for this flux suppression was being applied in Redrock during redshift fitting, leading to an over-suppression of template flux relative to what was expected at $\lambda_{\text{RF}}<1215.67$Å. This resulted in a bias that increases with the amount of suppression (i.e. redshift) at $z>2$.
@@
We use the updated HIZ templates with modified Redrock to refine the redshifts of the DESI year 1 quasar sample at $z>1.6$. This redshift is defined by the upper limit of the redshift coverage for the LOZ templates, though we note substantial impact is only expected at $z>1.95$. We impose a prior of $\Delta v=\pm 3000$ km s-1 from the original redshift estimate, as the HIZ redshifts have been shown to be approximately correct within a few tens to hundreds of km s-1, on average. Approximately 850,000 objects had a change in redshift when using the updated HIZ templates, and we show the distribution of those changes in figure 8.
Figure 8: Distribution of the difference in redshift (in km s-1) for the quasars in the year 1 catalog between the redshifts measured with the updated HIZ templates ($z_{\texttt{HIZ},\tau}$) versus the redshifts measured from the templates without the Lyman-$\alpha$ optical depth correction ($z_{\texttt{HIZ}}$).
@@
Parameter baseline z1 z2 z3 z4 No. quasars 510,501 201,290 138,364 82,710 88,137 Redshift range - $z\leq 2.2$ $2.2<z\leq 2.5$ $2.5<z\leq 2.8$ $z>2.8$ $b_{\text{Ly}\alpha}$ $-0.0543\pm 0.0035$ $-0.049\pm 0.005$ $-0.055\pm 0.005$ $-0.060\pm 0.006$ $-0.162\pm 0.020$ $\beta_{\text{Ly}\alpha}$ $3.78\pm 0.26$ $3.3\pm 0.4$ $3.8\pm 0.4$ $4.5\pm 0.5$ $2.01\pm 0.28$ $10^{3}b_{\eta,\text{SiII(1190)}}$ $-2.5\pm 0.4$ $-2.6\pm 0.4$ $-2.8\pm 0.4$ $-3.9\pm 0.5$ $-3.4\pm 0.5$ $10^{3}b_{\eta,\text{SiII(1193)}}$ $-1.34\pm 0.35$ $-1.3\pm 0.4$ $-2.2\pm 0.4$ $-2.0\pm 0.5$ $-2.2\pm 0.5$ $10^{3}b_{\eta,\text{SiIII(1207)}}$ $-4.50\pm 0.31$ $-5.0\pm 0.4$ $-5.0\pm 0.4$ $-6.0\pm 0.5$ $-5.4\pm 0.5$ $10^{3}b_{\eta,\text{SiII(1260)}}$ $-3.1\pm 0.6$ $-2.9\pm 0.5$ $-3.2\pm 0.6$ $-3.8\pm 0.7$ $-3.2\pm 0.6$ $b_{\text{HCD}}$ $-0.0574\pm 0.0034$ $-0.0491\pm 0.0029$ $-0.061\pm 0.004$ $-0.082\pm 0.005$ $-0.089\pm 0.006$ $b_{\text{q}}$ $3.97\pm 0.10$ $3.35\pm 0.13$ $3.99\pm 0.13$ $4.73\pm 0.17$ $4.35\pm 0.23$ $\Delta r_{\parallel}$ ($h^{-1}$pc) $-0.08\pm 0.04$ $-0.01\pm 0.06$ $-0.16\pm 0.07$ $-0.09\pm 0.08$ $-0.11\pm 0.13$ $\sigma_{\text{v}}$ ($h^{-1}$pc) $5.21\pm 0.17$ $4.79\pm 0.25$ $5.35\pm 0.25$ $5.45\pm 0.35$ $3.46\pm 0.35$ $\xi_{0}^{\text{TP}}$ $0.074\pm 0.008$ $0.066\pm 0.013$ $0.061\pm 0.014$ $0.114\pm 0.023$ $0.52\pm 0.09$ $10^{4}A_{\text{inst}}$ $1.33\pm 0.09$ $1.4\pm 0.1$ $1.36\pm 0.10$ $1.35\pm 0.10$ $1.51\pm 0.10$ $\chi^{2}$ 11941.2 11920.1 11883.1 11624.9 11573.8 $\chi^{2}$/dof 1.027 1.025 1.022 0.999 0.995 $z_{\text{eff}}$ 2.32 2.09 2.33 2.61 2.98
Table 4: Results of the fits using the updated quasar templates for the new baseline and four redshift bins. The number of quasars in each bin is given, along with the redshift range of each bin where applicable. The free parameters are the same as described in section 5. In all fits the BAO parameters are fixed at 1.0. Each fit has 11,644 data bins, 13 free parameters and we also list the effective redshift, $z_{\text{eff}}$. Since each fit is a joint fit of the auto- and cross-correlations, we list the $\chi^{2}$ and reduced $\chi^{2}$ ($\chi^{2}$/dof) values given for the joint fit.